From mar@roe.ac.uk Mon Oct 22 15:23:07 2012 Date: Mon, 22 Oct 2012 15:11:18 +0100 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Andrew Lawrence , Bob Mann , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Norman Gray , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS science archives project team meeting, 19 Oct 2012 Present: MAR, NCH, RSC, RPB, MSH Apologies: ETWS, NJC, RGM, JPE, AL, STV, NG, KTN DoNM: 10am Friday November 2nd 2012 *** NB in the Villas meeting room *** NCH in the chair. Actions discharged this time: ----------------------------- ACTION: NCH to check the GCS for pawsky-related issues, and liaise with SJW over the LAS results Discharged, NCH had written SQL to identify frames using pawskys based on a single position. Steve Warren wanted all the affected LAS frames deprecated. Phil Lucas and NCH to decide which of the GPS and GCS to deprecate. Actions partly discharged but continuing: ----------------------------------------- ACTION: ETWS (with help from NJC as necessary) to add FeII 1.644 micron filter details into the WSA. Continues; all done apart from Vega-to-AB and zeropoint attributes (the later needs the latest data to be ingested). Actions carried forward from 5/10/12 meeting --------------------------------------------- ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues; this one still has low priority. Specific points and new actions: -------------------------------- Project management: Nothing of note. WFCAM, VISTA and VST updates: Nothing to report this time. Comments and issues arising from CASU minutes: Nothing of note. Networking: The network was up and running following the power outage however the SDSS BestDR9 transfer had been cut short, ETWS to resume it on his return. We had received word from CASU that more VISTA 1.2 data are available for transfer. WSA/VSA/OSA/GES Operations: RPB reported that VVVDR1 and VSAVVV had been backed up. He'd also backed up the back up database! Fallout from the power outage had delayed much progress being made but all systems were back online and replication would be restored shortly allowing curation tasks to re-start. The disk38 copy was still proceeding. Hardware and Systems: It was thought that the issue with corrupted files on disk38 was due to the RAID controllers but investigations continue. Survey Data Release: With the pawsky issue in hand and SJW's LAS eyeball results in, UKIDSS DR10 can proceed through QC. VISTA releases are waiting on the disk38 copy to finish. With the VVVDR1 now backed up, the data originating from the bad Ks deep stacks could now be flagged/removed. ACTION: MAR to fettle VVVDR1 so that rows/columns in vvvSource and vvvDetection, originating from bad Ks deep stacks, are reset to defaults. Software: RSC reported working on CU6 speed ups. Non-survey Data Release: A new release had been made for WSERV2. Astrogrid deployment, VO & Data Analysis services: MSH reported that as of next week all surviving AG services would be hosted by WFAU. Miscellaneous: NJC had submitted an abstract for the ROE workshop. He had also been invited to give a talk at the "VISTA: a celebration" meeting at the ROE (17-18th January 2013). Others were encouraged to attend both meetings. MAR noted that Henrik Beuther of MPIA had been supplied with a bulk download script to retrieve GPS K-band pixel data. -- Scanned by iCritical.